OpenStudy (anonymous):

HELP PLZ AND FAST!! MEDAL AND FAN For which operations is the set {0, 1} closed? Choose all answers that are correct. A. multiplication B. division C. addition D. subtraction

OpenStudy (anonymous):

It is not subtraction.

OpenStudy (anonymous):

It would also be closed under multiplication

OpenStudy (anonymous):

so probably A and C

OpenStudy (anonymous):

ok thx

OpenStudy (anonymous):

can you help out with four more??

jimthompson5910 (jim_thompson5910):

It is NOT closed under addition because we can take the elements 1 and 1 and combine them to get 1+1 = 2 2 is NOT in the original set. So we can see that the set is NOT closed under addition

OpenStudy (anonymous):

oh ok

OpenStudy (anonymous):

But there is not two 1s in the set so it would seem not to be a valid argument but I could be wrong.

OpenStudy (anonymous):

ok i am so confused

OpenStudy (anonymous):

but i think i agree with bananaroof01

OpenStudy (anonymous):

but can you help me out with four more??

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

thx

OpenStudy (anonymous):

Which equations show that the set of whole numbers is not closed under subtraction? Choose all answers that are correct. A. 1 – (–2) = 3 B. 1 – 2 = –1 C. 2 – 0 = 2 D. 2 – 4 = –2

OpenStudy (anonymous):

ok so do you no how to do any of these qustions if not i can explan

OpenStudy (anonymous):

i do its just a bit confusing

OpenStudy (anonymous):

OK i will explan mid telling me what grad your in

jimthompson5910 (jim_thompson5910):

@bananaroof01 you can take any two elements of the set that you want. They can be repeated. The only rule is that the numbers picked must come from the set.

OpenStudy (anonymous):

im in 7th

OpenStudy (anonymous):

cool

OpenStudy (anonymous):

ok a is not going to be it and B and c would be whigt si do you want me to explane why

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

ok so A would not be it becuse it would not make sense

OpenStudy (anonymous):

are you there

OpenStudy (anonymous):

ya

OpenStudy (anonymous):

ok i just cretted my self it just c it maks the most sense

OpenStudy (anonymous):

do you have to pic to

OpenStudy (anonymous):

ya

OpenStudy (anonymous):

or more

OpenStudy (anonymous):

ok so i pic c and a

OpenStudy (anonymous):

fan me

OpenStudy (anonymous):

ok done

OpenStudy (anonymous):

did you get it right

OpenStudy (anonymous):

i still got like 3 more

OpenStudy (anonymous):

ok whats the nest one and do you go to k12

OpenStudy (anonymous):

yep

OpenStudy (anonymous):

me to

OpenStudy (anonymous):

cool

OpenStudy (anonymous):

Which sets of numbers are closed under addition? Choose all answers that are correct. A. {0, 2, 5, 8} B. even integers C. rational numbers D. {0}

OpenStudy (anonymous):

ok so we dont get band open a new post ok

OpenStudy (anonymous):

and tag me

OpenStudy (anonymous):

ok